OpenStudy (anonymous):

Which is the side length of a cube with a volume of 64 m3

OpenStudy (anonymous):

i mean what

OpenStudy (anonymous):

let side length of cube =x its volume will be=x^3 x^3=64 x=64^(1/3) x=4

OpenStudy (anonymous):

@SilverOwl

OpenStudy (anonymous):

thank you :)